kittybasil:

WORD PROBLEM: "One-third of all native flowering plant species in the United States are at risk of becoming extinct. That translates into 5144 at-risk flowering plant species. Based on this data, how many flowering plant species are native to the United States overall?" (Hint: How many \(\frac{1}{3}\)s are in 5144?) NOTE: I know how to solve this level of math. The issue is that the phrasing doesn't make sense... the Hint contradicts the question

Vocaloid:

The hint is kind of dumb, but if you think about how the problem is setup If x = all native plant species then (1/3)x = 5144 represents the proportion of at risk species When you do the calculations, you’ll, of course, divide 5144 / (1/3) which is equivalent to asking “how many 1/3’s are there in 5144”

kittybasil:

Oh, that's what they meant. ngl what I did was just read it as "well 1 out of 3 partitions is 5144, which means the total \(\frac{3}{3}\) is 3 \(\frac{1}{3}\)'s and thus 3 of \(5144\)" So... basically the same idea but I took a non-expression approach xD